OpenStudy (anonymous):

Before even beginning to solve the problem below, how do you know that 2 cannot be a solution of the equation? 3/(m - 2) + 1/(m - 1) = 1/ ( m^2 - 3m + 2)

OpenStudy (campbell_st):

2 in the 1st fraction will give a value of zero in the denominator, which is undefinded.

OpenStudy (campbell_st):

oops undefined

OpenStudy (radar):

same for 2nd equation, and 1 is a no no for 2nd equation, do you have to solve it?

OpenStudy (anonymous):

yes I need help solving it also.

OpenStudy (campbell_st):

factorise the right hand side \[\frac{3}{m-2} +\frac{1}{m -1} = \frac{1}{(m-2)(m-1)}\] multiply every term by (m-2)(m-1) which will give 3(m -1) +1(m -2) = 1 expand and collect like terms then solve for m

OpenStudy (radar):

3 1 1 -------- + ----------- = ---------------- m-2 m-1 (m-2)(m-1) factored (m^2 -3m +2) Multiply all terms by (m-2)(m-1) 3(m-1) + (m-2) = 1 3m-3+m-2=1 4m=4 m=1,,,,,,,but if m=1 then that means division by 0 so what are you to do!lol note the 2nd fraction would have a 0 denominator.